# Chinese vs. World Bank Development Projects: Insights from Earth Observation and Computer Vision on Wealth Gains in Africa, 2002-2013

### v1

@@ Abstract
Debates about whether development projects improve living conditions persist, partly because observational estimates can be biased by incomplete adjustment and because reliable outcome data are scarce at the neighborhood level. We address both issues in a continent‑scale, sector‑specific evaluation of Chinese and World Bank projects across 9,899 neighborhoods in 36 African countries (2002–2013), representative of $\sim$88% of the population. First, we use a recent dataset that measures living conditions with a machine‑learned wealth index derived from contemporaneous satellite imagery, yielding a consistent panel of 6.7 km square mosaics. Second, to strengthen identification, we proxy officials’ map‑based placement criteria using pre‑treatment daytime satellite images and fuse these with rich tabular covariates to estimate funder‑ and sector‑specific ATEs via inverse‑probability weighting. Incorporating imagery systematically shrinks effects relative to tabular‑only models, indicating prior work likely overstated benefits. On average, both donors raise wealth, with larger gains for China; sector extremes in our sample include Trade & Tourism for the World Bank (+6.27 IWI points), and Emergency Response for China (+14.32). Assignment‑mechanism analyses show World Bank placement is generally more predictable from imagery alone, as well as from tabular covariates. This suggests that Chinese project placements are more driven by non-visible, political, or event-driven factors than World Bank placements. To probe residual concerns about selection on observables, we also estimate within‑neighborhood (unit) fixed‑effects models at a spatial resolution about 450 times finer than prior fixed effects analyses, leveraging the computer-vision–imputed IWI panels; these deliver smaller but directionally consistent effects.
Keywords: Living conditions; World Bank programs; Chinese infrastructure programs; causal inference; Earth observation; computer vision; machine learning.
@@ Introduction
This study contributes to the literature by applying two new innovations that address data limitations and methodological challenges of evaluating the effectiveness of foreign aid from two distinct aid donors: China and the World Bank. We study the aid they provided to 36 African countries (representing 88% of the continent’s population) between 2002 and 2013, based on a representative sample of 9,899 neighborhoods as the pool of potential treatment and control units of analysis.
@@
We provide our EO-ML causal method with 5 km square daytime satellite images centered over treated and control communities, one period before the aid project commitment, along with a large set of tabular-format variables representing political, economic, and other factors that empirical research has shown are relevant to aid allocation and effectiveness. We employ EO-ML methods to identify features visible in satellite imagery (along with tabular covariates) to estimate each community’s treatment propensity. These propensities are used in an inverse probability weighting procedure to estimate the effect of aid from each funder and sector on neighborhood wealth.
This EO-ML approach targets the average treatment effect (ATE) estimand for 36 African countries over the period 2002 to 2013. Many policy evaluation studies on Chinese and World Bank programs target the local average treatment effect (LATE) estimand, because they use an instrumental variable approach. Although LATE is considered to provide a more credible identification approach, its estimand may struggle with external validity because the variation it targets is much narrower than the ATE, as this variation relies on the compliers in the instrument. As our goal is to estimate an Africa-wide effect over the given time period, we target the ATE. Our expectation is that we will improve our identification by relying on an additional adjustment set—satellite images to capture selection-relevant covariates.
@@ 1.1 The Impact of Chinese Development Finance on Living Conditions
China also funds numerous projects in social sectors, but the impact of these projects on living conditions is less studied and potentially more mixed. China’s demand-driven, state-led model may enable the rapid construction of schools and clinics, but some research suggests that these projects may be less effective at improving long-term outcomes compared to those of traditional donors (Martina, 2020). The reason often cited is a focus on “hardware” over “software”–Chinese projects tend to emphasize the provision of physical infrastructure without an equivalent investment in the staffing, training, and institutional capacity needed to deliver services effectively.
@@ 2 Two Key Gaps
Second is the persistent challenge of causal identification. Development programs are not randomly assigned Bedi et al. (2007b). While past studies have used instrumental variables or fixed-effects models, these approaches may not fully account for fine-grained confounders related to geography and the built environment. Our study contributes a new methodological approach to mitigate this problem: we use the satellite images themselves as control variables, which also enables new kinds of substantive insights to be gained about the assignment mechanism surrounding aid (Rubin, 1991; Daoud and Dubhashi, 2023)—capturing what kinds of information do and do not predict aid decisions. By conditioning on satellite images—which proxy for the maps used by aid officials in project allocation (Bedi et al., 2007a)—we compare treated and untreated areas with greater sensitivity, thereby enhancing the credibility of our causal estimates and addressing limitations in prior literature (Easterly, 2006). In this way, we can more robustly identify causal impacts of aid interventions while controlling for and explicitly characterizing the funder targeting strategies.
@@ 2.1 Measuring Strategic Allocation with Satellite Images & Computer Vision
Our first contribution addresses causal identification by accounting for the processes that determine treatment assignment in observational studies of development programs. We estimate the effects of World Bank and Chinese projects on neighborhood-level living conditions in Africa, drawing on the potential outcomes framework to define estimands and invoking graphical models to clarify assumptions. Our strategy leverages satellite imagery not merely as a covariate but as a proxy for the geospatial features that likely inform donors’ assignment decisions, thereby illuminating the program allocation mechanism itself Bedi et al. (2007b).
First, a note on the unit of analysis. As shown in Figure 1, the outcome $Y_{it}$ is a raster square of 6.7 km side lengths. The outcome represents socioeconomic well-being (i.e., health and living conditions broadly defined) at time $t$, measured via the International Wealth Index derived from satellite imagery as detailed in Pettersson et al. (2023) and further discussed in the data section. Substantively, this square $i$ measures approximately a neighborhood—a block in an urban area or a village in a rural area. The outcome is measured one 3-year-period after the treatment has been deployed, for simplicity, denoted $t+1$. The treatment $A_{it}$ denotes the presence or absence of a development project.
Each neighborhood has geocoordinates $\mathbf{l}_{i}\in\mathbb{R}^{2}$ (centroid of the square) and binary treatment indicators $A_{i,k,WB}\in\{0,1\}$ for World Bank projects and $A_{i,k,China}\in\{0,1\}$ for Chinese projects, where $k$ indexes sector-specific interventions (indices for actors and sectors are dropped for brevity where unambiguous). Sector definitions appear in the data section. A unit is in the treated group if a neighborhood $i$ meets at least one of the three criteria: the unit resides within a treated ADM2 boundary, has proximity within a 25 km radius of a project, or contains a project within its 6.7 km square boundary. Treatment status is determined contemporaneously with the commitment period.
@@
Identification commonly proceeds via backdoor adjustment, as depicted by the directed acyclic graph (DAG) in Figure 2(b). Confounding emerges when unobserved factors $U_{it}$ influence both assignment $A_{it}$ and outcome $Y_{it}$. In aid allocation, such $U_{it}$ often encompasses geospatial attributes—terrain texture, infrastructure layouts, proximity to natural features, and other strategic reasons—that donors consult via maps during site selection. These maps, unavailable to researchers, encode spatiotemporal selection logics that drive why certain neighborhoods receive aid over others. Satellite images $\mathbf{M}_{i(\overline{t}-1)}$, which capture both observable and latent geospatial traits (e.g., inferred historical condition on the ground), serve as proxies for these maps. Conditioning on $\mathbf{M}_{i(\overline{t}-1)}$ thus blocks confounding paths by approximating the assignment mechanism donors employ. The image $\mathbf{M}_{i(\overline{t}-1)}$, as discussed in Data, is measured as a composition of the previous three-year images, prior to the treatment (hence, the bar $\bar{t}$ over the time indicator in Figure 1).
@@
Adjusting for high-dimensional images $\mathbf{M}_{i(\bar{t}-1)}$ demands specialized methods. We employ a computer vision method to yield $f_{U}(\mathbf{M}_{i(\bar{t}-1)})$, integrated with $\mathbf{X}_{i(t-1)}$ for propensity estimation and inverse probability weighting. This not only facilitates ATE recovery but elucidates assignment by revealing geospatial drivers of aid placement. This approach is further detailed in Methods.
@@
To select neighborhoods, we use the locations of survey clusters from a round of the DHS for each country, as they are a nationally representative random sample of census enumeration areas. They approximate the size of a rural community or urban neighborhood, selected within urban and rural strata (Croft et al., 2018). At the time surveys were conducted, each cluster’s GPS coordinates were collected, and the coordinates were displaced for privacy purposes (Burgert et al., 2013). Excluding 11 DHS cluster locations that lacked adequate data, such as those with duplicate latitude and longitude coordinates or no population density estimates, we were left with 9,899 units of analysis. More details and descriptive statistics are provided in Tables DA1 and DA2 in the Data Appendix; see Table 1 for illustration.
@@ 3.1 Treatments: World Bank and China Development Projects
Having discussed our outcome data in the previous section, we now discuss the treatment. Data on both World Bank and China aid projects were sourced from (Dreher et al., 2021) and curated by AidData. Table 2 shows descriptive statistics, Figure 3 illustrates annual project counts, and Figure 4 depicts the sector distribution for the two funders over the study period. [^1] Aid projects spanning multiple sectors are treated by allocating the full project count to each relevant sector or subsector, without weighting or fractional division based on the number of sectors involved. The general takeaway is that the World Bank allocates more projects than China, and that Chinese projects primarily focus on health, transportation, and communication.
@@
Portion lacking funding information 100% 28%
@@
As discussed previously and shown in Figure 1, to define treatment and control groups and create panel data for each funder and sector, we define a neighborhood unit as part of the treatment group if any of the following three conditions is met: a precision 1 (exact) project lies within the 6.7 km neighborhood boundary, the neighborhood lies within a 25 km buffer around a precision 2 (near) project boundary, or the neighborhood lies within an all-ADM2 (precision 3) project area. If none of these conditions are fulfilled, the unit is assigned to the control group. To avoid confounding at the country selection level, we exclude neighborhoods from the control group in countries where the funder did not conduct any projects in the sector of interest during the study period.
@@
We exclude funder sector combinations with fewer than 100 treated locations, since they lack sufficient data for training and test splits. For the remaining sector/funder combinations, we check for and remove variables that lack variation across the treatment and control groups, as they add little statistical value (Table DA 3).
@@ 3.2 Tabular-Covariates Adjustment Set
We provide a large set of tabular confounders that the literature suggests influence both aid placement and effectiveness, variables for year fixed effects to account for common shocks, and ADM2 variables for fixed effects to account for numerous time-invariant factors at that level, including regional history, ethnic composition, and culture. There may be additional unobserved confounders that threaten inference. To avoid introducing post-treatment bias, all covariates are measured pre-treatment, ensuring they function as good controls (see Cinelli et al. (2024) for a detailed discussion). The operationalization of tabular variables and the methods used for their measurement are summarized in Table LABEL:tab:Descriptives and the section below. The Data Appendix contains more details.
@@
$\cdot$ Control of corruption score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Government effectiveness score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Political stability and absence of violence/terrorism score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Regulatory quality score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Rule of law score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Voice and accountability score from annual World Governance Indicators Country 3‑year average, one period prior to project commitment
@@
$\cdot$ Nightlights, per capita, logged 5 km 3‑year average, one period prior to project commitment
$\cdot$ Population density, average of 1 km squares within neighborhood, logged 5 km 3‑year average, one period prior to project commitment
@@
$\cdot$ Travel minutes to city of over 50,000 population in the year 2000, average of $\sim$1 km squares in neighborhood, logged 5 km Time invariant
@@
Because previous literature suggests both World Bank and China place projects in the non-poorest sub-national locations (Briggs, 2021; Dreher, 2019) and that projects are most successful in areas with brighter night lights (conditional on population) and above-average education (Briggs, 2021; Greßer and Stadelmann, 2021), we calculate a per-capita nightlight variable to represent a pre-project neighborhood wealth confounder. We use nightlight data from the Defense Meteorological Program Operational Linescan System (DMSP) and population data from Columbia University’s WorldPop Global Project (WorldPop, 2023). For each three-year period, we average the annual per capita nightlights and then add one, taking the log because the data is highly right-skewed.
Note that we are using past nightlight measurements, before the treatment was assigned, as a control variable. This will guard against post-treatment bias. Additionally, even if Pettersson et al’s method includes nightlight (and daylight jointly) to measure the IWI outcome variable, including nightlights again as a control will not bias our estimates. Essentially, by including historical nightlight levels as part of the confounding-adjustment set, we control for historical levels of socioeconomic trends that drive the treatment assignment.
@@
Since empirical research suggests both China and the World Bank allocate aid to more populous subnational regions (Dreher, 2019; Dreher et al., 2022a; Öhler et al., 2019) and population density can facilitate economic opportunity and market access (Scott, 2009), we create an annual average population density variable for the 5 km square area around each unit of observation neighborhood from Columbia University’s WorldPop Population Density dataset (WorldPop, 2023). Since the average population density is highly skewed, we add one and take the log after averaging the value over 3-year periods.
@@
Because empirical studies of both World Bank and China aid in Africa suggest funders avoid areas with high travel times to a city (Dreher et al., 2022a), and access to an urban area can provide economic opportunities that increase aid effectiveness (Scott, 2009), we calculate each neighborhood’s travel minutes to a city. To operationalize this, we use data on the estimated travel time to the nearest city of 50,000 or more inhabitants in the year 2000 (Weiss et al., 2018) to calculate the travel time in minutes for each point within the 5 km square area surrounding each unit of observation. We take the average for the square area, add one, and take the log.
@@
Because previous research suggests China gives more aid to provinces in years when they are the birthplace of the country’s leaders than in other years (Dreher, 2019), we use tabular, geocoded data on the birth region of political leaders and the years they were in power available from Harvard’s Political Leader’s Affiliation Database (PLAD) (Bomprezzi et al., 2024). We set the variable to 1 if the national executive leader in power during the annual or 3-year period was born in the neighborhood’s ADM1 and to 0 otherwise.
@@
Country-level World Governance Indicators.
A large body of literature argues that aid allocation and effectiveness depend on the quality of recipient country institutions (see Asatullaeva (2021) for a review). To capture multiple dimensions of recipient country governance and institutions, we use six country-level, annual variables from the World Bank’s World Governance Indicators (Kaufmann and Kraay, 2024): Control of Corruption, Government Effectiveness, Political Stability and Absence of Violence/Terrorism, Regulatory Quality, Rule of Law, and Voice and Accountability. Each neighborhood inherits the average score of its country each period, which is roughly normally distributed between -2.5 and 2.5.

### v2

@@ Abstract
Debates about whether development projects improve living conditions persist, partly because observational estimates can be biased by incomplete adjustment and because reliable outcome data are scarce at the neighborhood level. We address both issues in a continent‑scale, sector‑specific evaluation of Chinese and World Bank projects across 9,899 neighborhoods in 36 African countries (2002-2013), representative of $\sim$88% of the population. First, we use a recent dataset that measures living conditions with a machine‑learned wealth index derived from contemporaneous satellite imagery, yielding a consistent panel of 6.7 km square mosaics. Second, to strengthen identification, we proxy officials’ map‑based placement criteria using pre‑treatment daytime satellite images and fuse these with tabular covariates to estimate funder‑ and sector‑specific ATEs via inverse‑probability weighting. Incorporating imagery often shrinks effects relative to tabular‑only models. On average, both donors raise wealth, with larger and more consistent gains for China; sector extremes in our sample include Trade and Tourism (330) for the World Bank (+12.29 IWI points), and Emergency Response (700) for China (+15.15). Assignment‑mechanism analyses also show World Bank placement is often more predictable from imagery alone (as well as from tabular covariates). This suggests that Chinese project placements are more driven by non-visible, political, or event-driven factors than World Bank placements. To probe residual concerns about selection on observables, we also estimate within‑neighborhood (unit) fixed‑effects models at a spatial resolution about 67 times finer than prior fixed-effects analyses, leveraging the computer-vision-imputed IWI panels; these deliver smaller but, for Chinese projects, directionally consistent effects. Methodologically, we extend recent EO–ML causal inference frameworks by fusing pre‑treatment satellite imagery with tabular covariates to estimate treatment propensities, and by systematically benchmarking image‑augmented estimators against tabular‑only and unit fixed‑effects designs using new assignment‑mechanism diagnostics. Empirically, we provide a continent‑wide, sector‑specific comparison of the neighborhood‑level wealth effects of Chinese and World Bank projects across 9,899 African neighborhoods.
@@ Introduction
This study contributes to the literature by applying two innovations that address data limitations and methodological challenges of evaluating the effectiveness of foreign aid from two distinct aid donors: China and the World Bank. We study the aid they provided to 36 African countries (representing 88% of the continent’s population) between 2002 and 2013, based on a representative sample of 9,899 neighborhoods as the pool of potential treatment and control units of analysis.
@@
We provide our EO-ML causal method with daytime satellite images centered over treated and control communities, one period before the aid project commitment, along with a large set of tabular-format variables representing political, economic, and other factors that empirical research has shown are relevant to aid allocation and effectiveness. We employ EO-ML methods to identify features visible in satellite imagery (along with tabular covariates) to estimate each community’s treatment propensity. These propensities are used in an inverse probability weighting procedure to estimate the effect of aid from each funder and sector on neighborhood wealth.
This EO-ML approach targets the average treatment effect (ATE) estimand for 36 African countries over the period 2002 to 2013. Many policy evaluation studies on Chinese and World Bank programs target the local average treatment effect (LATE) estimand, because they use an instrumental variable approach. Although LATE is considered to provide a more credible identification approach, this estimand may struggle with external validity because the variation it targets is much narrower than the ATE, as this variation relies on the compliers in the instrument. As our goal is to estimate an Africa-wide effect over the given time period, we target the ATE. Our expectation is that we will improve our identification by relying on an additional adjustment set—satellite images to capture selection-relevant covariates.
@@ 1.1 The Impact of Chinese Development Finance on Living Conditions
China also funds numerous projects in social sectors, but the impact of these projects on living conditions is less studied and potentially more mixed. China’s demand-driven, state-led model may enable the rapid construction of schools and clinics, but some research suggests that these projects may be less effective at improving long-term outcomes compared to those of traditional donors (Martina, 2020). The reason often cited is a focus on “hardware” over “software”—Chinese projects tend to emphasize the provision of physical infrastructure without an equivalent investment in the staffing, training, and institutional capacity needed to deliver services effectively.
@@ 2 Two Key Gaps
Second is the persistent challenge of causal identification. Development programs are not randomly assigned Bedi et al. (2007b). While past studies have used instrumental variables or fixed-effects models, these approaches may not fully account for fine-grained confounders related to geography and the built environment. Our study adapts and applies an EO–ML causal adjustment approach to further enhance identification (Sakamoto et al., 2025): we use the satellite images themselves as control variables, which also enables new kinds of substantive insights to be gained about the assignment mechanism surrounding aid (Rubin, 1991; Daoud and Dubhashi, 2023)—capturing what kinds of information do and do not predict aid decisions. By conditioning on satellite images—which proxy for the maps used by aid officials in project allocation (Bedi et al., 2007a)—we compare treated and untreated areas with greater sensitivity, thereby enhancing the credibility of our causal estimates and addressing limitations in prior literature (Easterly, 2006). In this way, we can more robustly identify causal impacts of aid interventions while controlling for and explicitly characterizing the funder targeting strategies.
@@ 2.1 Measuring Strategic Allocation with Satellite Images & Computer Vision
To work towards closing these two gaps, we proceed as follows. We begin with identification: we formalize the project placement problem and show how pre-treatment satellite imagery can proxy the map-based information donors use when deciding where to locate projects, strengthening causal adjustment in an observational setting. We then turn to measurement, introducing the satellite-derived wealth panel that makes neighborhood-level evaluation feasible at continental scale.
We start by laying out the treatment-assignment problem and clarifying the role of imagery as a proxy for otherwise-unobserved spatial selection factors in observational studies of development programs. We will estimate the effects of World Bank and Chinese projects on neighborhood-level living conditions in Africa, drawing on the potential outcomes framework to define estimands and invoking graphical models to clarify assumptions. Our strategy leverages satellite imagery not merely as a covariate but as a proxy for the geospatial features that likely inform donors’ assignment decisions, thereby clarifying aspects of the program allocation mechanism itself Bedi et al. (2007b).
Here, a note on the unit of analysis is warranted. As shown in Figure 1, the outcome $Y_{it}$ is a raster of 6.7 km side lengths. The outcome represents socioeconomic well-being (i.e., health and living conditions broadly defined) at time $t$, measured via the International Wealth Index derived from satellite imagery as detailed in Pettersson et al. (2023) and further discussed in the data section. Substantively, this square $i$ measures approximately a neighborhood—a block in an urban area or a village in a rural area. The outcome is measured one 3-year-period after the treatment has been deployed, for simplicity, denoted $t+1$. The treatment $A_{it}$ denotes the presence or absence of a development project.
Each neighborhood has geocoordinates $\mathbf{l}_{i}\in\mathbb{R}^{2}$ (centroid of the square) and binary treatment indicators $A_{i,k,WB}\in\{0,1\}$ for World Bank projects and $A_{i,k,China}\in\{0,1\}$ for Chinese projects, where $k$ indexes sector-specific interventions (indices for actors and sectors are dropped for brevity where unambiguous). Sector definitions appear in the data section.
A unit is in the treated group if at least one of the following holds for projects in its vicinity: (i) for precision‑1 (exact) locations, the project point lies inside the 6.7 km neighborhood square to capture direct local exposure; (ii) for precision‑2 (“near”) locations, the neighborhood centroid is within 25 km of the reported project point to account for greater spatial uncertainty; or (iii) for precision‑3 locations, where AidData only reports the second administrative division (ADM2), the neighborhood falls inside that ADM2. In other words, we only treat entire ADM2s when project locations are intrinsically coarse (precision 3); for projects with more precise coordinates, we do not treat ADM2s as a whole. Treatment status is determined contemporaneously with the commitment period. [^1] Investigators who wish to apply alternative treatment-footprint rules (e.g., different buffer radii or point-in-polygon criteria) and assess the sensitivity of estimates can do so using our replication package.
@@
Identification commonly proceeds via backdoor adjustment, as depicted by the directed acyclic graph (DAG) in Figure 2(b). Confounding emerges when unobserved factors $U_{i(t-1)}$ influence both assignment $A_{it}$ and outcome $Y_{it}$. In aid allocation, such $U_{i(t-1)}$ often encompasses geospatial attributes—terrain texture, infrastructure layouts, proximity to natural features, and other strategic reasons—that donors consult via maps during site selection. These maps, unavailable to researchers, encode spatiotemporal selection logics that drive why certain neighborhoods receive aid over others. Satellite images $\mathbf{M}_{i(\overline{t}-1)}$, which capture both observable and latent geospatial traits (e.g., inferred historical condition on the ground), serve as proxies for these maps. Conditioning on $\mathbf{M}_{i(\overline{t}-1)}$ thus blocks confounding paths by approximating the assignment mechanism donors employ. The image $\mathbf{M}_{i(\overline{t}-1)}$, as discussed in Data, is measured as a composition of the previous three-year images, prior to the treatment (hence, the bar $\overline{t}$ over the time indicator in Figure 1).
@@
Adjusting for high-dimensional images $\mathbf{M}_{i(\overline{t}-1)}$ demands specialized methods. We employ a computer vision method to yield $f_{U}(\mathbf{M}_{i(\overline{t}-1)})$, integrated with $\mathbf{X}_{i(t-1)}$ for propensity estimation and inverse probability weighting. This not only facilitates ATE recovery but elucidates assignment by revealing geospatial drivers of aid placement. This approach is further detailed in Methods.
@@
We will later benchmark image‑based adjustment against simpler specifications (difference‑in‑means, tabular‑only, and ADM2 fixed effects) as well as unit-level fixed effects.
@@
To select neighborhoods, we use the locations of survey clusters from a round of the DHS for each country, as they are a nationally representative random sample of census enumeration areas. They approximate the size of a rural community or urban neighborhood, selected within urban and rural strata (Croft et al., 2018). At the time surveys were conducted, each cluster’s GPS coordinates were collected and then randomly displaced for privacy purposes—by up to 2 km in urban areas and 5 km in rural areas, with an additional 1% of rural clusters displaced up to 10 km (Burgert et al., 2013). Excluding 11 DHS cluster locations that lacked adequate data, such as those with duplicate latitude and longitude coordinates or no population density estimates, we were left with 9,899 units of analysis. More details and descriptive statistics are provided in Tables DA1 and DA2 in the Data Appendix; see Table 1 for illustration.
Our baseline outcome definition links each DHS cluster to the corresponding 6.7 km $\times$ 6.7 km IWI tile. This resolution is broadly comparable to the DHS GPS displacement used for confidentiality (up to 2 km in urban areas and up to 5 km in rural areas), but it does not guarantee that the undisplaced cluster location falls within the same tile. To assess sensitivity to this potential spatial misalignment, we also re-estimate our main models using a coarser outcome that averages the four adjacent IWI tiles (an effective 13.4 km $\times$ 13.4 km neighborhood; see Section 5.4).
@@ 3.1 Treatments: World Bank and China Development Projects
Having discussed our outcome data in the previous section, we now discuss the treatment. Data on both World Bank and China aid projects were sourced from (Dreher et al., 2021) and curated by AidData. Table 2 shows descriptive statistics, Figure 3 illustrates annual project counts, and Figure 4 depicts the sector distribution for the two funders over the study period. [^2] Aid projects spanning multiple sectors are treated by allocating the full project count to each relevant sector or subsector, without weighting or fractional division based on the number of sectors involved. The general takeaway is that China funds more projects overall (Table 2), but the World Bank’s portfolio spans many more distinct project types and locations.
@@
Portion lacking full funding information 100% 28%
@@
As discussed previously and shown in Figure 1, to define treatment and control groups and create panel data for each funder and sector, we define a neighborhood unit as part of the treatment group if any of the following three conditions is met: a precision 1 (exact) project lies within the 6.7 km neighborhood boundary, the neighborhood lies within a 25 km buffer around a precision 2 (near) project boundary, or the neighborhood lies within an all-ADM2 (precision 3) project area. [^3] Note that the ADM2‑based rule applies only to precision‑3 locations; when AidData reports precision‑1 or precision‑2 coordinates, we rely solely on the 6.7 km and 25 km square rules and do not treat entire ADM2s. If none of these conditions are fulfilled, the unit is assigned to the control group. To avoid confounding at the country selection level, we exclude neighborhoods from the control group in countries where the funder did not conduct any projects in the sector of interest during the study period.
@@
We exclude funder-sector combinations with fewer than 80 treated locations, as they lack sufficient data for training and test splits. For the remaining sector/funder combinations, we check for and remove variables that lack variation across the treatment and control groups, as they add little statistical value.
@@ 3.2 Tabular-Covariates Adjustment Set
We provide a large set of tabular confounders that the literature suggests influence both aid placement and effectiveness, variables for year fixed effects to account for common shocks, and ADM2 variables for fixed effects to account for numerous time-invariant factors at that level, including regional history, ethnic composition, and culture. There may be additional unobserved confounders that threaten inference. To avoid introducing post-treatment bias, covariates are measured pre-treatment where possible, making it more likely that they function as good controls (see Cinelli et al. (2024) for a detailed discussion). Tabular variable measurement is summarized in Table LABEL:tab:Descriptives and the section below. The Data Appendix contains more details.
@@
$\cdot$ Control of corruption score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Government effectiveness score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Political stability and absence of violence/terrorism score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Regulatory quality score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Rule of law score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
$\cdot$ Voice and accountability score from annual Worldwide Governance Indicators Country 3‑year average, one period prior to project commitment
@@
$\cdot$ Nightlights, per capita, logged 6.7 km 3‑year average, one period prior to project commitment
$\cdot$ Population density, average of 1 km squares within neighborhood, logged 6.7 km 3‑year average, one period prior to project commitment
@@
$\cdot$ Travel minutes to city of over 50,000 population in the year 2000, average of $\sim$1 km squares in neighborhood, logged 6.7 km Time invariant
@@
Because previous literature suggests both World Bank and China place projects in the non-poorest sub-national locations (Briggs, 2021; Dreher, 2019) and that projects are most successful in areas with brighter night lights (conditional on population) and above-average education (Briggs, 2021; Greßer and Stadelmann, 2021), we calculate a per-capita nightlight variable to represent a pre-project neighborhood wealth confounder. We use nightlight data from the Defense Meteorological Satellite Program Operational Linescan System (DMSP) and population data from Columbia University’s WorldPop Global Project (WorldPop, 2023). For each three-year period, we average the annual per capita nightlights and then add one, taking the log because the data are highly right-skewed.
Because nightlights are persistent and also enter the Pettersson et al. (2023) outcome-construction pipeline, we also conduct robustness checks that drop lagged per-capita nightlights from the adjustment set. Results are reported in Section 5.4.
@@
Since empirical research suggests both China and the World Bank allocate aid to more populous subnational regions (Dreher, 2019; Dreher et al., 2022a; Öhler et al., 2019) and population density can facilitate economic opportunity and market access (Scott, 2009), we create an annual average population density variable for the square areas around each unit of observation neighborhood from the WorldPop Population Density dataset (WorldPop, 2023). Since the average population density is highly skewed, we add one and take the log after averaging the value over 3-year periods.
@@
Because empirical studies of both World Bank and China aid in Africa suggest funders avoid areas with high travel times to a city (Dreher et al., 2022a), and access to an urban area can provide economic opportunities that increase aid effectiveness (Scott, 2009), we calculate each neighborhood’s travel minutes to a city. To operationalize this, we use data on the estimated travel time to the nearest city of 50,000 or more inhabitants in the year 2000 (Weiss et al., 2018) to calculate the travel time in minutes for each point within the 6.7 km square area surrounding each unit of observation. We take the average for the square area, add one, and take the log.
@@
Because previous research suggests China gives more aid to provinces in years when they are the birthplace of the country’s leaders than in other years (Dreher, 2019), we use tabular, geocoded data on the birth region of political leaders and the years they were in power available from the Political Leaders’ Affiliation Database (PLAD) (Bomprezzi et al., 2024). We set the variable to 1 if the national executive leader in power during the annual or 3-year period was born in the neighborhood’s ADM1 and to 0 otherwise.
@@
Country-level Governance Indicators.
A large body of literature argues that aid allocation and effectiveness depend on the quality of recipient country institutions (see Asatullaeva (2021) for a review). To capture multiple dimensions of recipient country governance and institutions, we use six country-level, annual variables from the World Bank’s Worldwide Governance Indicators (Kaufmann and Kraay, 2024): Control of Corruption, Government Effectiveness, Political Stability and Absence of Violence/Terrorism, Regulatory Quality, Rule of Law, and Voice and Accountability. Each neighborhood inherits the average score of its country each period, which is roughly normally distributed between -2.5 and 2.5.